OpenStudy (anonymous):

Please help to evaluate the integral: ∫x^2 ln (x^2)dx

OpenStudy (anonymous):

you would have to use integration by parts.....

OpenStudy (australopithecus):

Thats what I thought lol

OpenStudy (anonymous):

then your thoughts were right....i am sure you were doubting yourself..

OpenStudy (australopithecus):

I haven't done an integral since last august lol

OpenStudy (anonymous):

Yes thanks, if I let u = x^2 then du = 2x dx and dv = ln (x^2), what would v = ?

OpenStudy (australopithecus):

The formula for integration by parts, integral f(x)g'(x)dx = f(x)g(x) - integral f'(x)g(x)dx

OpenStudy (anonymous):

hope this link helps. http://www.youtube.com/watch?v=dqaDSlYdRcs...

OpenStudy (australopithecus):

There is a really good video on khan academy but I cant seem to find it :(

OpenStudy (anonymous):

http://www.youtube.com/watch?v=g-Uxw0eKuJo is that it??

OpenStudy (australopithecus):

Maybe, Anyways, Take your integral and split it into two functions ∫x^2 ln (x^2)dx for example set, f(x) = x^2 g'(x) = ln(x^2) then input them in the formula f(x)g(x) - integral f'(x)g(x)dx You should try to do this in a way so that you get the simplest integral possible I hope this explanation is helpful but you may want to check out the posted videos it is easier when someone explains it to you

OpenStudy (anonymous):

Thanks Australopithecus and kausarsalley - I understand the formula, but am struggling with the bit where I need to find the integral of ln (x^2). Appreciate your help!

OpenStudy (australopithecus):

or does solves a problem with it rather

OpenStudy (australopithecus):

REMEMBER LOG RULES! ln(x^2) = 2ln(x)

OpenStudy (australopithecus):

I suspect you can solve it now ;)

OpenStudy (australopithecus):

You should have a table that tells you the derivative of ln(x)

OpenStudy (australopithecus):

oh sorry lol, been awhile since i have done an integral, I'm pretty sure that is something you are suppose to memorize

OpenStudy (australopithecus):

you can just use u substitution I'm pretty sure though

OpenStudy (australopithecus):

set u = ln(x)

OpenStudy (anonymous):

Yes I need to memorize it indeed! derivative of ln(x) is 1/x, but would integral of 2lnx then be 2xlnx - 2x? I think I'll try u = ln(x^2) yes thanks!

OpenStudy (australopithecus):

looks good to me

OpenStudy (anonymous):

Thanks for your help - I ended up using u = ln(x^2) then du = 2/x dx and dv = x^2 then v = x^3/3 and using the formula managed to get the answer.